# CorMulT: A Semi-supervised Modality Correlation-aware Multimodal Transformer for Sentiment Analysis

### v1

@@ 4.6. Ablation Studies
To clearly verify the effectiveness of our MCE modality-correlated pre-training model and the internal details of MCE, we conducted two sets of ablation studies. The studies respectively validate the influence of MCE multimodal feature extraction modules and the number of pre-trained modalities incorporated into the CorMulT model.
In our Modality-Specific Feature Extraction module, we provided specialized treatment for the uniqueness of the three types of data. To validate its impact, we specifically analyze the Mel spectrum matrix transformation for audio data and the ResNet50 image feature mapping.
Table 5. The ablation study results for MCE multimodal feature extraction.
Strategy L_A L_T L_V L_ave Epoch_time
@@
From the results, the specialized treatment strategies in the Modality-Specific Feature Extraction module showed significant effectiveness. By using both strategies, LOSS_ave was reduced from 0.315 to 0.168, and the training time decreased from 178.24 to 41.02.
We also tested our CorMulT model under various pre-trained modality combinations, including pre: comb, pre: A+T, pre: V+T, and pre: V+A+T.
Table 6. The ablation study results of pre-trained modality quantity.
Modality Acc@7 Acc@2 F1 MAE Corr
V+A+T(pre: comb) 50.9 82.9 83.4 0.553 0.752
V+A+T(pre: A+T) 51.2 83.2 83.8 0.554 0.762
V+A+T(pre: V+T) 52.3 86.2 86.1 0.552 0.778
V+A+T(pre: V+A+T) 52.8 86.5 86.7 0.552 0.779
In summary, the performance of the CorMulT model improves as more modalities are considered during pre-training, indicating that inter-modality information is a significant factor in enhancing sentiment analysis performance. Compared to other models, CorMulT displays superior stability and performance due to its effective use of modality correlations.
5. Conclusion
The main contributions and innovations of this paper can be summarized in three aspects. Firstly, a novel multimodal correlation pre-training model called MCE is proposed, which effectively extracts features from three modalities: audio, image, and text, and maps them into a shared joint space to learn the correlations between the modalities. Secondly, a semi-supervised contrastive learning training architecture is designed, which generates correlation data using different strategies without the need for manually annotated datasets, thereby improving the efficiency and feasibility of the model pre-training. Lastly, by combining MCE with the state-of-the-art multimodal fusion method MULT, the best performance of multimodal correlation pre-training is achieved, surpassing other existing research methods.
The experimental results demonstrate that CorMulT achieves an F1 score of 86.7% on the CMU_MOSEI dataset, which is an improvement of over 3.2% compared to the baseline model and outperforms the existing MMLatch model, making it the best-performing model among all multimodal sentiment analysis models. The ablation experiments further confirm the importance of the multi-modal correlation model in sentiment analysis models. Overall, the experimental results strongly support the significance of mutual information between modalities in improving the performance of sentiment analysis models. With its excellent stability and powerful performance, the CorMulT model demonstrates the effective utilization of intermodal correlation and provides a new solution for the field of multimodal sentiment analysis.
Acknowledgements.
This work was supported in part by Open Research Projects of Zhejiang Lab (No. 2021KH0AB04).

### v2

@@ IV-F Ablation Studies
We conducted two sets of ablation studies to validate the influence of MCE multimodal feature extraction module and the number of modalities used for pre-training CorMulT model.
For modality-specific feature extraction, we applied tailored approaches to address the distinct characteristics of three different data types. To gauge the effectiveness of these approaches, we specifically examined the transformation of the Mel spectrum matrix for audio data and the application of ResNet50 for image feature mapping. Table VI quantitatively evaluates the effectiveness of different feature extraction strategies in the MCE module across various modalities. The strategies include the use of Mel spectrum transformation for audio data, ResNet50 for image data, both methods combined, and a baseline with no specialized feature extraction methods. Metrics assessed include loss values for audio ($Loss_{A}$), text ($Loss_{T}$), and visual ($Loss_{V}$) data, along with the average loss ($Loss_{ave}$) and epoch time ($t_{epoch}$). The results demonstrate the efficacy of modality-specific feature extraction. Tailored approaches such as Mel spectrum for audio and ResNet50 for visual data significantly reduce the losses across respective modalities, with the combined approach yielding the lowest average loss ($Loss_{ave}$ = 0.168) and shortest training time (41.02 seconds per epoch). This verifies the importance of employing specific feature extraction methods for each modality.
TABLE V: The results of CorMulT while using different modality-specific feature extraction methods.
Strategy $Loss_{A}$ $Loss_{T}$ $Loss_{V}$ $Loss_{ave}$ $t_{epoch}$
@@
We also assessed CorMulT’s performance with different combinations of pre-trained modalities, including no modality specified (None), audio and text only (A+T), visual and text only (V+T), and visual, audio, and text (V+A+T). The results shown in Table VI demonstrate that CorMulT’s performance is enhanced by incorporating more modalities during pre-training. The results underscore the significance of inter-modality synergy in improving sentiment analysis outcomes. This matches the expectation that the more modality similarity information that CorMulT learns, the higher performance it achieves.
TABLE VI: The results of CorMulT while using different modalities for pre-training.
Modalities Acc@7 Acc@2 F1 Corr
None 50.9 82.9 83.4 0.752
A+T 51.2 83.2 83.8 0.762
V+T 52.3 86.2 86.1 0.778
V+A+T 52.8 86.5 86.7 0.779
V Conclusion
In this work, we proposed the two-stage semi-supervised model CorMulT which designs a modality correlation contrastive learning module to efficiently learn modality correlation coefficients between different modalities, and fuses the learned correlation coefficients with modality representations to make accurate sentiment prediction. According to the experimental evaluation, CorMulT obviously outperform existing multimodal sentiment analysis methods. In addition, the ablation studies also demonstrate the significance of mutual information between modalities in improving the performance of sentiment analysis models. With its excellent stability and high performance, CorMulT model achieves effective utilization of intermodal correlations and provides a new solution to multimodal sentiment analysis. As for future work, it is promising to further explore the dynamics of intermodal correlations for better multimodal sentiment analsyis.